# AI Rules? Characterizing Reddit Community Policies Towards AI-Generated Content

### v1

@@ Abstract.
How are Reddit communities responding to AI-generated content? We explored this question through a large-scale analysis of subreddit community rules and their change over time. We collected the metadata and community rules for over $300,000$ public subreddits and measured the prevalence of rules governing AI. We labeled subreddits and AI rules according to existing taxonomies from the HCI literature and a new taxonomy we developed specific to AI rules. While rules about AI are still relatively uncommon, the number of subreddits with these rules almost doubled over the course of a year. AI rules are also more common in larger subreddits and communities focused on art or celebrity topics, and less common in those focused on social support. These rules often focus on AI images and evoke, as justification, concerns about quality and authenticity. Overall, our findings illustrate the emergence of varied concerns about AI, in different community contexts.
@@ 1. Introduction
Of course, generative AI and AIGC may impact the rules (Makyen, 2023), dynamics (Hua et al., 2024; Fiesler, 2024), and governance (Cao et al., 2023) strategies of online communities in complex ways. While some recent research has explored how generative AI might help online communities deal with problematic content (Agarwal et al., 2024; Kumar et al., 2024), other work has documented the ability of AIGC to pollute these same communities with deceptive information (Park et al., 2024). Recent work by Lloyd et al. (Lloyd et al., 2023) used interviews with Reddit moderators to explore the particular challenges of governing AI use in online communities and found that communities are enacting explicit rules about AIGC in order to clarify norms. Lloyd et al.’s work focuses on the perspectives and experiences of their interviewees and raises the question of how the range of concerns about AIGC that the moderators voice, both practical and ideological, generalize across the wide variety of communities on Reddit.
This paper addresses that question through a large-scale quantitative analysis of community rules on Reddit that govern AI usage. Our approach allows us to paint a fuller picture for the HCI community by complementing Lloyd et al.’s qualitative findings with an analysis of platform-wide patterns and trends, while simultaneously providing deeper insight into the specific types of AI rules that communities are creating. We build on quantitative techniques developed by Fiesler et al. (2018) to study the general ecosystem of community rules on Reddit and adapt their methodology to specifically investigate the emergence of community rules that govern AI use. We use the same community topic and rule taxonomies as Fiesler et al. so that we can compare our analysis of AI rules to their analysis of Reddit rules in general. We address the following research questions:
  • RQ1: How common are subreddit rules governing the use of AI and how has this changed over the last year?
@@
To answer these questions we crawled over $300,000$ public subreddits twice, during the summers of 2023 and 2024, in order to build a dataset of $n=80,290$ English language subreddits with rules. Our inquiry into RQ1 finds that while only $3\%$ of subreddits with rules have rules about AI, the number of subreddits with AI rules has almost doubled between the time of our crawls. More specifically, larger subreddits are much more likely to have AI rules: for example, of the top $1\%$ largest subreddits with rules, almost $19\%$ had AI rules in our most recent crawl. To explore RQ2, we perform subreddit-level analysis on all subreddits in our sample with rules. We apply a “Topic” label to these subreddits using a taxonomy from Fiesler et al. (Fiesler et al., 2018) and “Community Archetype” labels using a taxonomy from Prinster et al. (Prinster et al., 2024). We use quantitative methods to explore the relationship between these labels and the presence of AI rules and find these rules to be most common in communities with Art and Celebrity topics, as well as those that fit the Content Generation Community Archetype. To explore RQ3, we perform rule-level analysis on all the rules in our sample that we identified as governing AI use. We apply an existing rule taxonomy from Fiesler et al. (Fiesler et al., 2018) and additionally produce our own set of emergent labels to describe meaningful dimensions of variation in both the requirements imposed by these rules and the language through which they frame AI usage. We find rules about AI images to be more common than rules about other types of AIGC and see quality and authenticity as the most frequently expressed concerns.
We compare these findings to the results of related work and discuss implications for platform designers. Our findings identify some common community concerns about quality and authenticity, especially related to images, that may be addressed through thoughtful design. Our findings also show, though, that communities of different types take varied approaches towards regulating AI use, and suggest that design solutions meant to help communities address these challenges must be sensitive to community context.
@@ 2.1. Online Communities’ Rules, Norms, and Values
Reddit is a fruitful platform for a study of community self-moderation because it is home to online communities of various size and purpose. This heterogeneity, combined with its largely public and text-based nature, have made Reddit a frequent topic of study in online communities research (Proferes et al., 2021). Recent work has explored the values (Weld et al., 2024), norms (Chandrasekharan et al., 2018), and rules (Fiesler et al., 2018) of different communities across the platform, both via direct methods like surveys, and indirect methods such as analyzing the digital trace data produced in these communities. Reddit’s implementation of community self-moderation allows communities to write and make public their own community-enforced rules, which Fiesler et al. collect and characterize in a study that is the main methodological inspiration for our work (Fiesler et al., 2018). In their study, Fiesler et al. find subreddit size to be the main predictor of having rules, but note that even among subreddits of similar sizes rules are context-dependent, not uniform, and often uniquely tailored to the topic of the subreddit (e.g., art subreddits have more rules about copyright). The authors also highlight that setting norms and developing rules is fundamentally a social process. Overall, Fiesler et al. show the importance of community rules and the feasibility of using them as a data source to study attitudes and stances across a platform. Given the potential impact of AI on online communities, we use community rules as a starting point to study context-specific attitudes towards AIGC as they emerge across the platform.
@@ 2.2. Generative AI’s Impact on Online Communities
Many additional studies have started to document the impacts of AIGC on online communities “in the wild”. Recent work has identified undisclosed AI use in several settings, such as LLM-powered social bots on Twitter (Yang and Menczer, 2024), AIGC in fake online reviews (Oak and Shafiq, 2024), and spam content used for scams (DiResta and Goldstein, 2024). In addition to studies of deceptive use, other recent work has begun to explore how the disclosed use of AIGC may affect community dynamics, such as in art-sharing communities (Wei and Tyson, 2024). Generative AI can additionally impact online communities by drawing traffic away from communities and decreasing overall engagement: causal inference studies of the Q&A sites in the Stack Exchange network, where users gather to ask and answer questions about programming and related topics, have demonstrated that the launch of ChatGPT caused an overall decline in website visits and question volume (Burtch et al., 2024). The authors of this study contrast this decline with a measured null-effect on similar Reddit communities, which they suggest may avoid a decline in traffic because of their emphasis on socialization, rather than pure information exchange. Finally, recent work by Lloyd et al. used interviews to qualitatively explore the attitudes and experiences of Reddit moderators dealing with AIGC (Lloyd et al., 2023). This study interviewed moderators from subreddits of different sizes and topics and surfaces a variety of community objections to AIGC that range from practical to ideological. The work poses open questions about how prevalent these concerns are across the platform and suggests that larger trends could be explored quantitatively in future work. Our study fills that gap through a quantitative analysis that helps identify the most common concerns and the most impacted communities, so that the HCI community can begin the task of addressing these issues with context-specific solutions.
@@ 3. Methods
In order to perform a large-scale analysis of community rules governing the use of AI, we built four datasets consisting of the metadata of public, English-language subreddits through a process summarized in Figure 1. These datasets of subreddits and their associated rules include data for the same subreddits collected at two different points in time, one year apart. The datasets include: the Broad Subreddit Set ($n=264,592$) consisting of subreddits with and without rules, the Rules Subreddit Set ($n=80,290$) of only subreddits with explicit rules, the AI Rules Subreddit Set ($n=2,600$) of subreddits with rules governing the use of AI, and the AI Rules Set ($n=4,150$) of individual rules governing AI use. We then classified the AI Rules Set using emergent coding, through which we created a taxonomy of AI rules. We supplement the raw crawl data with additional information about subreddit and rule types through several rounds of labeling, using both human and LLM coders. We will make our datasets and labels public upon completion of the peer-review process.
@@
A funnel diagram with five steps showing the filtering of subreddits based on various criteria: 1. Step 1: 337,399 subreddits, representing 100% of all subreddits seen during the 2023 crawl. 2. Step 2: 295,152 subreddits (87.5%), representing subreddits still accessible during the 2024 crawl. 3. Step 3: 264,592 subreddits (78.4%), filtered to only subreddits where English is the primary language. 4. Step 4: 80,290 subreddits (23.8%), filtered to subreddits with non-empty ”rules” during either the 2023 or 2024 crawl. 5. Step 5: 2,600 subreddits (0.771%), further filtered to subreddits with rules governing AI use during either crawl. The final output is 4,150 individual AI rules identified in these subreddits.
3.1. The Broad and Rules Subreddit Sets
@@
To build the Broad Subreddit Set of $n=264,592$ public, English-language subreddits, we performed two automated crawls: one in the summer of 2023 and another one year later, in summer 2024. The first crawl collected metadata for all subreddits indexed by Reddit’s “Top Communities” list [^1] https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We used a web-scraping library [^2] https://docs.scrapy.org/en/latest/ to visit the webpage for each of these subreddits and extract relevant metadata from the page HTML. This first crawl scraped $n=337,399$ subreddits in total.
@@
We performed our second crawl in the summer of 2024. We fetched up-to-date metadata for each subreddit seen during our first crawl. As our main goal here was understanding rule development over time, we removed from our sample $n=42,247$ subreddits that were inaccessible during this second crawl [^3] Subreddits could be inaccessible either because they no longer existed or were not public at the time that we attempted to crawl them.. We used an automated tool [^4] https://pypi.org/project/langdetect/ to determine the language of each subreddit based on the text in its description and rules metadata. We removed $n=30,560$ subreddits from our sample whose primary language was not English. This process left us with the $n=264,592$ of the Broad Subreddit Set which we used to understand platform-wide patterns in the data and factors that contribute to whether subreddits have rules.
@@
Using this broad dataset, we constructed a second, smaller dataset: the Rules Subreddit Set ($n=80,290$). This dataset consists only of subreddits with non-empty rules metadata at the time of at least one of our crawls. We use this dataset to understand patterns specific to subreddits with rules. These two datasets are the focus of our descriptive analysis and were used to create two other datasets specific to AI rules, which are discussed in Sections 3.2.2 and 3.2.3.
@@ 3.2. Data Labeling
To answer our research questions about rules governing the use of AI, we needed labels that identified these rules and provided additional context about both the rules and the communities that used them. Through the process described below, we added Topic, Community Archetype and Has AI Rule? labels to each of the subreddits in the Rules Subreddit Set. From these labels we identified the subset of subreddits that had AI rules, which we refer to as the AI Rules Subreddit Set ($n=2,600$). The specific AI-related rules within these subreddits became our AI Rules Set ($n=4,150$) and were each labeled with Rule Type, AI Rule: Rationale, AI Rule: Stance, AI Rule: Medium, and AI Rule: Framing labels. Below, we describe the labels for subreddits and rules and the process through which we created them. The applied labels are summarized in Table 1 (subreddits) and Table 2 (rules).
Table 1. Labels applied to subreddits in the Rules Subreddit Set
Label Source Mutually Exclusive? Example Values
Topic Fiesler et al. Yes Art, Celebrity, Video Games
Community Archetype Prinster et al. No
Content Generation, Topical Q&A,
Social Support
Has AI Rule? New Yes Yes, No
@@
Label Source Mutually Exclusive? Example Values
@@ 3.2.1. Detecting Community Rules Governing AI Use
Community rules are often very long and complex, which made the Rules Subreddit Set too large to manually code for the presence of rules governing the use of AI. To get a rough sense of the prevalence of these rules, we first used a simple regex to search the rules in the dataset for AI-related keywords. This step found matches in the rules of less than $2\%$ of the subreddits. We then devised a protocol, inspired by recent HCI work using LLMs for qualitative labeling tasks (Choksi et al., 2024; Rao et al., 2024; Aubin Le Quéré et al., 2024), to detect AI-related rules across the entire dataset with more precision. The approach is based on comparing LLM-produced labels with those produced by human experts and iterating on the LLM prompt until desirable inter-rater reliability (IRR) is achieved.
@@
We used these manually labeled rules as ground truth against which to evaluate the performance of an LLM coder, which we planned to use to label the rest of the dataset. We crafted a prompt that combined the human coder instructions with Chain-of-Thought prompting techniques (Wei et al., 2023), including the recommended eight few-shot, in-context examples, which we selected from the manually labeled sample. We produced LLM labels by using OpenAI’s API to prompt gpt-4o-mini, OpenAI’s current flagship small model, using this prompt and the same input data that were given to the human coders (a subreddit’s rules). We compared the LLM-produced labels with the human-produced ones and calculated an IRR of .7.7.7.7, which we deemed insufficient. To improve IRR we repeated this process of randomly sampling, manually coding, generating LLM labels, comparing the results, and updating the prompt using the examples that the model got wrong. We performed three iterations of this step, at which point the LLM labels had an IRR of .95.95.95.95 with the human labels. We then labeled the entire Rules Subreddit Set using OpenAI’s API and gpt-4o-mini. The final prompt used can be found in Appendix A.1. This process identified the AI Rules Subreddit Set ($n=2,600$) of subreddits with rules governing AI use.

### v2

@@ Abstract.
How are Reddit communities responding to AI-generated content? We explored this question through a large-scale analysis of subreddit community rules and their change over time. We collected the metadata and community rules for over $300,000$ public subreddits and measured the prevalence of rules governing AI. We labeled subreddits and AI rules according to existing taxonomies from the HCI literature and a new taxonomy we developed specific to AI rules. While rules about AI are still relatively uncommon, the number of subreddits with these rules more than doubled over the course of a year. AI rules are more common in larger subreddits and communities focused on art or celebrity topics, and less common in those focused on social support. These rules often focus on AI images and evoke, as justification, concerns about quality and authenticity. Overall, our findings illustrate the emergence of varied concerns about AI, in different community contexts. Platform designers and HCI researchers should heed these concerns if they hope to encourage community self-determination in the age of generative AI. We make our datasets public to enable future large-scale studies of community self-governance.
@@ 1. Introduction
Of course, generative AI and AIGC may impact the rules (Makyen, 2023), dynamics (Hua et al., 2024; Fiesler, 2024), and governance (Cao et al., 2023) strategies of online communities in complex ways. While some recent research has explored how generative AI might help online communities deal with problematic content (Agarwal et al., 2024; Kumar et al., 2024), other work has documented the ability of AIGC to pollute these same communities with deceptive information (Park et al., 2024). Recent work by Lloyd et al. (Lloyd et al., 2023) used interviews with Reddit moderators to explore the particular challenges of governing AI use in online communities and found that communities are enacting explicit rules about AIGC in order to clarify norms. Lloyd et al.’s work focuses on the perspectives and experiences of their interviewees and raises the question of how their findings generalize across the wide variety of communities on Reddit.
This paper addresses that question through a large-scale quantitative analysis of community rules on Reddit. Our approach allows us to paint a fuller picture for the HCI community by complementing Lloyd et al.’s qualitative findings with an analysis of platform-wide patterns and trends, while simultaneously providing deeper insight into the specific types of AI rules that communities are creating. We build on techniques developed by Fiesler et al. (2018) to study the ecosystem of community rules on Reddit and adapt their methodology to specifically investigate the emergence of community rules governing AI use. We use Fiesler et al.’s community topic and rule taxonomies so that we can compare our analysis of AI rules to their analysis of general Reddit rules. We address the following research questions:
  • RQ1: How common are subreddit rules governing the use of AI and how has their prevalence changed over time?
@@
To answer these questions we crawled over $300,000$ public subreddits on two occasions, in July 2023 and November 2024, building a dataset of $N=99,969$ English-language subreddits with published rules. Analyzing a longitudinal sample of subreddits that were present during both our crawls, our inquiry into RQ1 finds that while only $1.2\%$ of subreddits now have rules about AI, the number of subreddits with AI rules has more than doubled since July 2023. Larger subreddits are much more likely to have AI rules: for example, of the top $1\%$ largest subreddits, $17\%$ had AI rules in our most recent crawl. To explore RQ2, we perform subreddit-level analysis on all subreddits in our 2024 sample with rules. We apply a “Topic” label to these subreddits using a taxonomy from Fiesler et al. (Fiesler et al., 2018) and “Community Archetype” labels using a taxonomy from Prinster et al. (Prinster et al., 2024). We use quantitative methods to explore the relationship between these labels and the presence of AI rules and find these rules to be most common in communities with Art and Celebrity topics, as well as those that fit the Content Generation Community Archetype. To explore RQ3, we perform rule-level analysis on the rules in our sample that we identified as governing AI use. We apply an existing rule taxonomy from Fiesler et al. (Fiesler et al., 2018) and additionally produce our own set of emergent labels to describe meaningful dimensions of variation in both the requirements imposed by these rules and the language through which they frame AI usage. We find rules about AI images to be more common than rules about other types of AIGC and see quality and authenticity as the most frequently expressed concerns.
We compare these findings to the results of related work and discuss implications for platform designers. In particular, thoughtful design choices may help address the community concerns about quality and authenticity that our findings identify. Our findings also show, though, that communities of different types take varied approaches towards regulating AI use, and suggest that design solutions meant to help communities address these challenges must be sensitive to community context.
As an additional contribution, we make our community rules data and labels public. While we limit our current analysis to questions about AI rules, our data can be used to answer a wide range of questions about content moderation in self-moderating online communities. We hope that this will enable large-scale analysis by future HCI researchers interested in encouraging online community self-determination.
@@ 2.1. Online Communities’ Rules, Norms, and Values
Reddit is a fruitful platform for a study of community self-moderation because it is home to online communities of various size and purpose. This heterogeneity, combined with its largely public and text-based nature, have made Reddit a frequent topic of study in online communities research (Proferes et al., 2021). Recent work has explored the values (Weld et al., 2024), norms (Chandrasekharan et al., 2018), and rules (Fiesler et al., 2018) of different communities across the platform, both via direct methods like surveys, and indirect methods such as analyzing digital trace data from these communities. Reddit’s implementation of community self-moderation allows communities to write and make public their own community-enforced rules, which Fiesler et al. collect and characterize in a study that is the main methodological inspiration for our work (Fiesler et al., 2018). In their study, Fiesler et al. find subreddit size to be the main predictor of having rules, but note that even among subreddits of similar sizes rules are context-dependent, not uniform, and often uniquely tailored to the topic of the subreddit (e.g., art subreddits have more rules about copyright). The authors also highlight that setting norms and developing rules is fundamentally a social process. Overall, Fiesler et al. show the importance of community rules and the feasibility of using them as a data source to study attitudes and stances across a platform. Given the potential impact of AI on online communities, we use community rules as a starting point to study context-specific attitudes towards AIGC as they emerge across Reddit.
@@ 2.2. Generative AI’s Impact on Online Communities
Many additional studies have started to document the impacts of AIGC on online communities “in the wild”. Recent work has identified undisclosed AI use in several settings, such as LLM-powered social bots on Twitter (Yang and Menczer, 2024), AIGC in fake online reviews (Oak and Shafiq, 2024), and spam content used for scams (DiResta and Goldstein, 2024). In addition to studies of deceptive use, other recent work has begun to explore how the disclosed use of AIGC may affect community dynamics, such as in art-sharing communities (Wei and Tyson, 2024). Generative AI can additionally impact online communities by drawing traffic away from communities and decreasing overall engagement: causal inference studies of Q&A sites in the Stack Exchange network, where users gather to ask and answer questions about programming and related topics, have demonstrated that the launch of ChatGPT caused an overall decline in website visits and question volume (Burtch et al., 2024). The authors of this study contrast this decline with a measured null-effect on similar Reddit communities, which they suggest may avoid a decline in traffic because of their emphasis on socialization, rather than pure information exchange. Finally, recent work by Lloyd et al. used interviews to qualitatively explore the attitudes and experiences of Reddit moderators dealing with AIGC (Lloyd et al., 2023). This study interviewed moderators from subreddits of different sizes and topics and surfaces a variety of community objections to AIGC that range from practical to ideological. The work poses open questions about how prevalent these concerns are across the platform and suggests that larger trends could be explored quantitatively in future work. Our study fills that gap through a quantitative analysis that helps identify the most common rules and the most impacted communities, so that the HCI community can begin the task of addressing these issues with context-specific solutions.
@@ 3. Methods
In order to perform a large-scale analysis of community rules governing the use of AI, we collected the metadata of public, English-language subreddits. Building on a large-scale crawl in 2024, and an earlier one performed in 2023, we created five datasets of subreddits and their associated rules. We processed the more recent crawl to create four datasets: the Broad Subreddit Set ($N=307,543$) consisting of the metadata of subreddits with and without rules; the Rules Subreddit Set ($N=99,969$) of only subreddits with stated, published rules; the AI Rules Subreddit Set ($N=4,251$) of subreddits with rules governing the use of AI; and the AI Rules Set ($N=4,458$) of individual rules governing AI use. The process of constructing these datasets is summarized in Figure 1, and detailed below. We supplement the raw crawl data with additional information about subreddit and rule types through several rounds of labeling, using both human and LLM coders. We use the AI Rules Set and emergent coding to construct a novel taxonomy of AI Rule Types. In addition, we created a longitudinal dataset of subreddits that were present in both the 2023 and 2024 crawls. We use this Longitudinal Subreddit Set ($N=227,737$) to compare the prevalence of AI rules at different points in time for a consistent sample of subreddits. We will make our datasets and labels public upon completion of peer-review.
@@
A funnel diagram with five steps showing the filtering of subreddits based on various criteria: 1. Step 1: 331,938 subreddits, representing 100% of all subreddits seen during the 2024 crawl. 2. Step 2: 307,543 subreddits (92.7%), filtered to subreddits older than 1 month old where English is the primary language. 3. Step 3: 99,969 subreddits (30.1%), filtered to subreddits with non-empty ”rules”. 4. Step 4: 4,251 subreddits (23.8%), further filtered to subreddits with rules governing AI use. The final output is 4,458 individual AI rules identified in these subreddits.
3.1. The Broad Subreddit and Rules Subreddit Sets
@@
In order to better understand how online communities across Reddit are responding to the arrival of AIGC, we conducted a large-scale data collection and analysis. In November 2024 we performed an automated crawl of Reddit’s “Top Communities” list [^1] https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We used a web-scraping library [^2] https://docs.scrapy.org/en/latest/ to paginate through this list and extract the IDs of all listed subreddits from the page HTML. We then used Reddit’s Data API [^3] https://support.reddithelp.com/hc/en-us/articles/16160319875092-Reddit-Data-API-Wiki to fetch metadata and rules for each subreddit ID.
@@
From the $N=331,938$ subreddits that we saw in our crawl, we used a process summarized in Figure 1 to create several datasets for analysis. We used an automated tool [^4] https://pypi.org/project/langdetect/ to determine the language of each subreddit based on the text in its description and rules metadata. We removed subreddits with a primary language other than english or a creation date less than one month before the crawl date, in order to remove communities that had not yet had time to create rules. This yielded the Broad Subreddit Set of $N=307,543$ public, English-language subreddits that were at least one month old, which we used to understand platform-wide patterns in the data and factors that contribute to whether subreddits have rules at all.
@@
Using this broad dataset, we constructed a second, smaller dataset: the Rules Subreddit Set ($N=99,969$). This dataset consists only of subreddits with non-empty rules metadata. We use this dataset to understand patterns specific to subreddits with rules. These two datasets are the focus of our descriptive analysis and were used to create two other datasets specific to AI rules, which are discussed in Sections 3.3.4 and 3.3.2.
3.2. Longitudinal Subreddit Set
To address RQ1’s concern with the shift in rule prevalence over time, we constructed a Longitudinal Subreddit Set, using a July 2023 data collection effort as a baseline, and identifying the subset of subreddits from the 2024 crawl that were seen in the earlier crawl. ChatGPT launched at the end of 2022, kicking off a period of increased public awareness and usage of AI. Given the dynamic nature of such a moment of flux, we decided to perform a longitudinal analysis on the same set of subreddits over a period of time, so that we could study the evolution of community norms and practices soon after this launch.
The 2023 crawl used the same technique as the 2024 crawl, capturing subreddits in Reddit’s “Top Communities” list as described above. In the 2023 crawl, we scraped the subreddit metadata and rules data directly from the web page of each subreddit, using the same tool that we used to visit the “Top Communities” list—a simpler approach that was possible under Reddit’s prior rate limit. We verified that the data that we gathered from a subreddit’s web page is identical to the data returned by the API, which we used for the 2024 crawl. This 2023 crawl scraped $N=337,399$ subreddits in total. From this 2023 dataset, we identified the subreddits whose primary language was English, and that were also present in the 2024 crawl. This process yielded the Longitudinal Subreddit Set ($N=227,737$) which we used to study the change in prevalence of AI rules over time.
@@ 3.3. Data Labeling
To answer our research questions about rules governing the use of AI, we created multiple label categories for the AI rules and the communities that had rules. Through the process described below, we used a hybrid human-LLM coding technique to determine whether each of the subreddits in the Rules Subreddit Set and the Longitudinal Subreddit Set had an explicit rule about AI use, and to label them accordingly. We used these labels to identify the subset of subreddits with AI rules. We refer to the subset of the Rules Subreddit Set with AI rules as the AI Rules Subreddit Set ($N=4,251$). The AI-related rules from this set became our AI Rules Set ($N=4,458$). We used an iterative inductive coding technique to identify common types of AI rules. We then used a hybrid human-LLM coding technique to label the rules according to these categories.
Below, we describe the label categories we used to categorize subreddit AI rules (summarized in Table 2), as well as the human-LLM process that we used to label the data. We also describe how we determined and labeled the community type for each subreddit in the Rules Subreddit Set, according to several frameworks from existing literature (summarized in Table 2).
@@
Label Category Source Mutually Exclusive? Example Labels
@@
Table 2. Labels applied to subreddits in the Rules Subreddit Set
Label Category Source Mutually Exclusive? Example Labels
Topic Fiesler et al. Yes Art, Celebrity, Video Games
Community Archetype Prinster et al. No Content Generation, Topical Q&A, Social Support
Has AI Rule? New Yes Yes, No
@@ 3.3.1. Detecting Community Rules Governing AI Use
Community rules are often very long and complex, which made our datasets too large to manually code for the presence of rules governing the use of AI. To get a rough sense of the prevalence of these rules, we first used a simple regex to search the subreddits’ rules for AI-related keywords. This step found matches in the rules of less than $2\%$ of the subreddits. We then devised a protocol, inspired by recent HCI work using LLMs for qualitative labeling tasks (Choksi et al., 2024; Rao et al., 2024; Aubin Le Quéré et al., 2024), to detect AI-related rules across the larger datasets with more precision. The approach is based on comparing LLM-produced labels with those produced by human experts and iterating on the LLM prompt until desirable inter-rater reliability (IRR) is achieved between the human and LLM labels.
@@
We used these manually labeled rules as ground truth against which to evaluate the performance of an LLM coder, which we planned to use to label the rest of the data. We crafted a prompt that combined the human coder instructions with Chain-of-Thought prompting techniques (Wei et al., 2023), including the recommended eight few-shot, in-context examples, which we selected from the manually labeled sample. We produced LLM labels by using OpenAI’s API to prompt gpt-4o-mini-2024-07-18, OpenAI’s current flagship small model, using this prompt and the same input data that were given to the human coders (a subreddit’s rules). We compared the LLM-produced labels with the human-produced ones and calculated an IRR of .7.7.7.7, which we deemed insufficient. Two human coders examined the discrepancies, discussed what they thought the correct label should be, and updated the prompt with a clarifying example. For example, one rule labeled NO by humans but YES by the LLM was about bots: “No bots(No bots): No bots. All bots are banned unless the bot provides exceptional value to /r/ MealPrepSunday.” To better align the LLM label with our preferred definition of an AI rule, we added this rule as a negative example in the prompt, with the explanation:“This rule bans bots, not AI. Bots are not necessarily AI.”
To improve IRR we repeated this process of randomly sampling, manually coding, generating LLM labels, comparing the results, and updating the prompt using the examples that the model got wrong. We performed three iterations of this step, at which point the LLM labels had an IRR of .95.95.95.95 with the human labels. We then labeled the entirety of both the Longitudinal Subreddit Set and the Rules Subreddit Set using OpenAI’s API and gpt-4o-mini-2024-07-18. The final prompt used can be found in Appendix A.1. This process identified the AI Rules Subreddit Set ($N=4,251$) of subreddits with rules governing AI use.
3.3.2. Identifying Emergent AI Rule Types
In order to better understand the types of rules that subreddits use to govern AI, we identified these rules in our sample and used an iterative inductive coding process to develop a descriptive taxonomy. To do this coding, we slightly modified the prompt developed and validated in Section 3.3.1 so that it would flag individual rules rather than a set of rules as being AI-related. Since the prompt required only a slight modification from the prior prompt, and since human coders would be manually reviewing the labels in the next step, we decided that we did not need to validate model performance before tagging the entire sample. We used this prompt to instruct gpt-4o-mini-2024-07-18 to label each of the $N=33,674$ individual rules in the AI Rules Subreddit Set, which produced the AI Rules Dataset ($N=4,458$) of individual rules about AI.
Next, two of the authors collaborated to identify emergent labels for the rules. The authors randomly selected $N=25$ rules from this dataset and independently performed affinity mapping (Harboe and Huang, 2015) by grouping rules that appeared similar into emergent categories. The authors then compared results and consolidated their categories into labels that best captured the elements of interest. Each of the two authors then independently applied this set of labels to another random sample of $N=25$ rules, once again grouping similar rules, including into new emergent groups where the existing labels did not suffice. The two authors repeated this process of randomly sampling, independently labeling, collaboratively comparing their results, and consolidating their labels, until no new labels emerged and they were confident that their taxonomy covered the relevant dimensions of variation in the dataset. The final label categories were Rationale, Stance, Medium and Framing, which are listed alongside some of their possible labels in Table 2 (see Table 6 for a full list of all labels and examples of tagged rules). We applied these labels to the entire AI Rules Dataset, as we describe next.
3.3.3. Applying Rule Classification Labels
We used LLMs to apply our emergent labels and labels from an additional Rule Type category from Fiesler et al. (Fiesler et al., 2018) to the entire AI Rules Dataset. The two first authors collaboratively developed a codebook with instructions for applying these labels. Fiesler et al. does not provide codebook instructions, so the authors interpreted the examples presented in the study and came up with instructions for how to apply each of the $24$ possible labels555“Rule Type” labels are not mutually exclusive, so we represented them as a series of binary labels, one for each possible Rule Type..
[…]